# Forecasting Communication Derailments Through Conversation Generation

### v1

@@ 6.3 Social Orientation Labels for Understanding Conversation Dynamics
Social orientation labels have been proposed to analyze the dynamics in conversations Morrill et al. (2024). Social orientation is a concept originally developed from Circumplex Theory Leach et al. (2015); Jonathan et al. (2005); Koch et al. (2016) in Psychology. The Circumplex Model allows for the analysis of interpersonal interactions along defined axes like power, benevolence, arousal, and political leaning, each with contrasting traits (e.g. assertive vs. submissive for power). In our proposed approach, we adopt social orientation annotations to guide our LLM in generating future exchanges.
6.4 Chain-of-Thought in LLMs
The Chain-of-Thought (CoT) method Wei et al. (2022) enhances an LLM’s reasoning abilities on complex tasks by generating intermediate steps before arriving at the final answer. This approach has shown effectiveness in tasks like mathematical problem-solving Wei et al. (2022); Wang et al. (2023); Fu et al. (2023), code generation Fu et al. (2023), sentiment analysis Fei et al. (2023); Sun et al. (2023); Yao et al. (2024). Chain-of-Thought with Self-Consistency Wang et al. (2023) further improves the original method by sampling multiple potential responses and selecting the most probable intermediate step through majority voting.
Our methodology could be viewed as a form of CoT in a broad sense. Unlike prior CoT approaches for sentiment analysis Fei et al. (2023); Sun et al. (2023); Yao et al. (2024), which generate step-by-step reasoning based on existing text, our method generates conversation continuations as intermediate steps. Similar to CoT, our approach allows LLMs to leverage test-time computation and solve tasks that are difficult to address directly.
@@ 7 Conclusion and Future Work
We introduced a novel approach for forecasting communication derailments by generating potential future conversation trajectories based on existing conversation history. This technique demonstrated strong performance on communication derailment prediction benchmarks, such as the CGA-Wiki and BNC datasets. We validated the effectiveness of our method through comparisons with state-of-the-art models and comprehensive ablation studies. We also assessed the contribution of social orientation labels in guiding derailment prediction. Future research could extend our methodology to conversations beyond online discourse, such as in-person conversations and meetings.
@@ 9 Ethical Considerations
For our study on predicting conversation derailment, we used two publicly available datasets: Conversation Gone Awry Wiki Split (CGA-Wiki), licensed under the MIT license, and Before Name Calling (BNC), licensed under Apache 2.0. We believe our use of these datasets complies with fair-use guidelines. Neither dataset contains personally identifiable information. However, because these datasets are intended to identify conversational derailments, they inevitably include offensive language.
@@ A.1 Conversation Generation Examples
A.2 Impact of Number of Generated Conversation Continuations
CGA-Wiki (Mistral + G) BNC (Mistral + SO + G)
# of Generations (L) Acc. Prec. Rec. F-1 Acc. Prec. Rec. F-1
1 68.5 66.0 76.2 70.7 92.7 92.4 93.1 92.7
3 70.2 67.1 79.3 72.7 93.1 91.8 94.6 93.2
5 71.4 68.4 79.5 73.6 93.8 93.2 94.6 93.9
Table 4: Accuracy, precision, recall, and F-1 scores of our best-performing model on CGA-Wiki and BNC, ablated by the number of generated future conversations per sample at test time. Scaling the number of generations consistently improves all metrics on both datasets.
We conducted an ablation study to examine the impact of $L$, the number of generated conversation continuations, on the performance of our best-performing methods across datasets, as shown in Table 4. Specifically, we evaluated each sample using 1, 3, and 5 continuations, applied majority voting, and measured accuracy, precision, recall, and F1-score. Due to computational resource constraints, we limited the maximum number of continuations per sample to 5. The results support our hypothesis that increasing $L$ improves the approximation of future conversational trajectories, leading to consistent performance gains across all metrics as number of continuations increases.

### v2

@@ 6.3 Social Orientation Labels for Understanding Conversation Dynamics
Social orientation labels are proposed to analyze the dynamics in conversations Morrill et al. (2024). Social orientation is a concept originally developed from Circumplex Theory Leach et al. (2015); Jonathan et al. (2005); Koch et al. (2016) in Psychology. The Circumplex Model allows for the analysis of interpersonal interactions along defined axes like power, benevolence, arousal, and political leaning, each with contrasting traits (e.g. assertive vs. submissive for power). In our proposed approach, we adopt social orientation annotations to guide our LLM in generating future exchanges.
@@ 7 Conclusion and Future Work
We introduced a novel approach for forecasting conversation derailments by generating potential future conversation trajectories based on existing conversation history. This technique demonstrated strong performance on conversation derailment prediction benchmarks, such as the CGA-Wiki and BNC datasets. We validated the effectiveness of our method through comparisons with state-of-the-art models and comprehensive ablation studies. We also assessed the contribution of social orientation labels in guiding derailment prediction. Future research could extend our methodology to conversations beyond online discourse, such as in-person conversations and meetings.
@@ 9 Ethical Considerations
For our study on predicting conversation derailment, we used two publicly available datasets: CGA-Wiki, licensed under the MIT license, and BNC, licensed under Apache 2.0. We believe our use of these datasets complies with fair-use guidelines. Neither dataset contains personally identifiable information. However, because these datasets are intended to identify conversational derailments, they inevitably include offensive language.
@@
10 Supplementary Materials Availability Statement
Our code, dataset, and model weights are available at this GitHub repository: https://github.com/YunfanZhang42/ConversationDerailments.
11 Acknowledgments
This research is being developed with funding from the Defense Advanced Research Projects Agency (DARPA) Cross-Cultural Understanding program under Contract No HR001122C0034. The views, opinions and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
@@ A.1 Conversation Generation Examples
A.2 Ablation Studies on Number of Continuations
CGA-Wiki (Mistral + SO + G) BNC (Mistral + SO + G)
# of Generations (L) Acc. Prec. Rec. F1 Acc. Prec. Rec. F1
1 64.2 64.7 62.7 63.7 89.2 84.9 94.4 89.4
3 64.7 65.0 63.9 64.4 89.9 86.7 93.6 90.0
5 66.5 66.9 65.3 66.1 90.3 86.8 94.4 90.4
7 64.9 65.6 62.9 64.2 90.3 86.2 95.2 90.5
11 65.3 66.0 63.2 64.5 90.7 86.9 95.2 90.8
15 66.6 67.6 63.9 65.7 90.3 86.2 95.2 90.5
Table 4: Accuracy, precision, recall, and F1 scores of the Mistral-based classifier on the validation sets of CGA-Wiki and BNC, varying the number of generated future conversations per sample at test time. SO denotes stands for orientation labels, and G stands for generation. Increasing the number of generations per conversation ($L$) from 1 to 5 improves performance on both datasets, but further increases beyond $L=5$ yield diminishing returns. We therefore set $L=5$ in all experiments to balance predictive performance with computational efficiency.
To determine $L$, the number of continuations per test sample, we conduct an ablation study on the CGA-Wiki and BNC validation sets, experimenting with $L$ between 1 and 15.
Intuitively, generating multiple conversation continuations and applying a majority vote over the resulting derailment predictions can reduce the variance introduced by the stochastic nature of LLM sampling, thereby improving prediction accuracy. However, generating too many continuations per conversation can be computationally costly.
Due to the large number of generations required for this experiment, we adopt vLLM Kwon et al. (2023) to improve generation speed, while the rest of the experiments in the paper use native PyTorch to ensure numerical consistency between training and testing environment.
As shown in Table 4, increasing the number of generations $L$ from 1 to 5 leads to 1-2% improvement in accuracy, along with improvement in precision, recall, and F1. Beyond $L=5$, however, additional gains are marginal. Based on these findings, we set $L=5$ for all experiments in this paper to maintain a reasonable trade-off between predictive accuracy and computational overhead.